# One Model to Rule them All: Towards Universal Segmentation for Medical Images with Text Prompts

### v2

@@ Abstract
Abstract. In this study, we focus on building up a model that aims to Segment Anything in medical scenarios, driven by Text prompts, termed as SAT. Our main contributions are three folds: (i) for dataset construction, we combine multiple knowledge sources to construct the first multi-modal knowledge tree on human anatomy, including 6502 anatomical terminologies; Then we build up the largest and most comprehensive segmentation dataset for training, by collecting over 22K 3D medical image scans from 72 segmentation datasets with careful standardization on both image scans and label space; (ii) for architecture design, we formulate a universal segmentation model, that can be prompted by inputting medical terminologies in text form. We present knowledge-enhanced representation learning on the combination of a large number of datasets; (iii) for model evaluation, we train a SAT-Pro with only 447M parameters, to segment 72 different segmentation datasets with text prompt, resulting in 497 classes. We have thoroughly evaluated the model from three aspects: averaged by body regions, averaged by classes, and average by datasets, demonstrating comparable performance to 72 specialist nnU-Nets, i.e., we train nnU-Net models on each dataset/subset, resulting in 72 nnU-Nets with around 2.2B parameters for the 72 datasets. We will release all the codes, and models in this work.
@@ 1 Introduction
Figure 1: Different from conventional specialist models (a) that develop specialized solution for each task, or recently proposed interactive universal segmentation models (b) relying on real-time human interventions, Segment Anything with Text (SAT) utilizes text prompts as queries and 3D volumes as inputs to perform a wide array of medical image segmentation tasks across different modalities, anatomies, and body regions (c). It can be easily applied to clinics or seamlessly integrated with any agent-based large language model.
@@
There is a growing interest in developing universal models for medical image segmentation [68, 37], by adapting the pre-trained Segment Anything Model (SAM) [46] models from the computer vision community. However, while transferring to medical scenarios, these models trained on natural images suffer from fundamental limitations: (i) models typically perform 2D slice segmentation, that are later fused into 3D volumes through post-processing. This approach overlooks the critical contextual information inherent in 3D radiological imaging; (ii) models often require point or box inputs as prompts, thus is an interactive segmentation model, requiring considerable manual efforts for use in practice; (iii) models suffer from significant domain gaps, from image statistics to domain-specific medical knowledge, e.g., ‘liver is distinctive for positioning in the upper-right part of the abdominal cavity, filling almost the entire right hypochondrium and often extending into the left hypochondrium as far as the mammillary line’.
@@
In this paper, we present the first knowledge-enhanced universal model for 3D medical volume segmentation, with language/text as prompt, termed as SAT. In practice, our model can effectively take 3D volume as visual inputs along with text prompts, to seamlessly tackle various medical image segmentation tasks, across modalities, anatomies, and body regions. As illustrated in Fig. 1, our proposed method distinguish from previous medical segmentation paradigms, can be seamlessly applied to clinics or integrated with any large language model. Specifically, we make the following contributions:
@@
On dataset construction, we construct a knowledge tree based on multiple medical knowledge sources, encompassing thousands of anatomy concepts and definitions throughout the human body. On the visual side, we curate over 22K 3D medical image scans with 302K anatomical segmentation annotations, covering 497 categories from 72 publicly available medical segmentation datasets, termed as SAT-DS. To the best of our knowledge, SAT-DS denotes the largest and most comprehensive collection of public 3D medical segmentation datasets. To achieve such goal, we have invested significant efforts in standardizing datasets and unifying annotation labels, paving the way for universal segmentation task. For complete list of datasets and download links, we refer the reader to Table LABEL:tab:dataset_details and Table LABEL:tab:dataset_links.
@@
Figure 2: Overview of our dataset for training universal segmentation model. Specifically, we collect datasets from a wide range of segmentation tasks across various modalities and anatomical regions of the human body, including the brain, head and neck, thorax, spine, abdomen, upper limb, lower limb, and pelvis.
@@
On architecture design and training strategy, we build a universal medical segmentation model that leverages text prompts to guide segmentation, that enables flexible segmentation across a spectrum of medical imaging modalities. Specifically, we adopt knowledge-enhanced representation learning, leveraging textual anatomical knowledge and atlas segmentation of specific anatomical structure to train the text encoder. Through this training process, the visual features of these anatomical structures are aligned with their corresponding text descriptions in the latent space. Later, the text embeddings of anatomy/abnormality are treated as queries in a Transformer-based architecture, enabling the model to selectively extract task-specific features for precise segmentation of the queried target. To satisfy the requirements from different computational resources, we train two models of varying sizes, namely, SAT-Nano, and SAT-Pro.
@@
On evaluation, we devise comprehensive metrics for universal medical segmentation across various aspects, including region-wise average, organ-wise average, and dataset-wise average. Through extensive experiments, we demonstrate that:
@@
  • Building on the unprecedented dataset collection, SAT-Pro with only 447M parameters, can employ flexible text prompts for a wide range of downstream segmentation tasks, showing comparable performance to specialist nnU-Net models trained individually on each dataset, i.e., results from 72 separate nnU-Net models.
  • Benefiting from the standardized data processing and unified label system, SAT show excellent generalisation ability when performing zero-shot transfer to clinic data. It is worth emphasizing that SAT can flexibly satisfy the clinical needs of doctors from different departments, without any additional annotation and training.
  • Enhanced by multimodal medical domain knowledge, our text encoder provides superior guidance for universal medical segmentation on 3D inputs, surpassing the state-of-the-art language model tailored for medical tasks. For the first time, we demonstrate how domain knowledge would facilitate the segmentation task.
  • While enlarging model capacity, i.e., from SAT-Nano to SAT-Pro, would significantly improve the segmentation performance. This reveals the promising yet still under-exploited potential of a universal medical segmentation model.
  • The text-prompted feature distinguishes SAT from all previous universal medical segmentation models that require human interaction. We demonstrate how SAT can be directly integrated with large language models like GPT-4 [2], automatically segmenting the targets of interests without human intervention.
@@ 2 Results
The goal of our proposed SAT is to build an universal segmentation for 3D medical images, driven by text prompt. Such method should be capable of handling diverse segmentation tasks excellently in one model, without laborious specialization and training for each single dataset. In addition, the universality should make it adaptable to clinic procedures with minimal extra efforts, addressing a broad range of clinical needs; and the text-driven character should equip it with immense potential as a flexible toolkit, bridging between language and segmentation. Also, to meet different hardware requirements, we train two variants of SAT, namely SAT-Pro and SAT-Nano.
2.1 Main Results
Building an universal segmentation model poses several new challenges compared to specialist models, such as the heterogeneity of data, the long-tail distribution, the limitation from model capacity and so on. To this end, we have constructed a large-scale training set, namely SAT-DS dataset, covering a wide spectrum of medical image segmentation tasks, covering 497 anatomical targets on 8 regions and lesions of the human body, across 72 datasets. We train two universal segmentation models, namely, SAT-Pro and SAT-Nano, and compare them with the powerful specialist model, i.e., nnU-Nets trained on each dataset. To comprehensively understand such capability, we conduct the evaluations from the perspective of anatomical regions, classes, and datasets respectively. Note that, class-wise and region-wise evaluations require averaging results across different nnU-Nets. For example, the results on ‘brainstem’ in CT images is the macro average of nnU-Nets trained on ‘HAN Seg’, ‘PDDCA’ and ’SegRap2023 Task 1’, since all these datasets contains annotation for this class.
2.1.1 Region-wise Results.
Table 1 shows the segmentation performance of our methods and nnU-Nets on 8 regions of human body, including Brain, Head and Neck, Thorax, Abdomen, Pelvis, Spine, Upper Limb, and Lower Limb, as well as Lesion, in terms of DSC and NSD, respectively. Classes existing in multiple regions are specifically grouped as Whole Body. Notably, SAT-Pro consistently outperforms nnU-Nets in three regions: on Head and Neck, it surpasses nnU-Nets by 5.21 on DSC and 8.27 on NSD; On Upper Limb, it surpasses nnU-Nets by 2.8 on DSC and 9.28 on NSD; On Lower Limb, it leads to 1.53 on DSC and 2.48 on NSD. SAT-Nano rank last in almost all the regions, except for Spine, where it slightly exceeds SAT-Pro by 0.82 on DSC and 0.81 on NSD.
For comparison, we visualize the results in Fig. 3 (c). Overall, SAT-Pro shows comparable segmentation performance to the 72 nnU-Nets, that are specialized for each dataset, and outperforms SAT-Nano in most cases. In addition, it’s worth noticing that SAT-Pro not only performs better, more flexible, but also significantly smaller, approximately 1/5 of the assemble of nnU-Nets; While SAT-Nano is even smaller, remarkably only 1/20 of the assemble of nnU-Nets in size.
Figure 3: Comparison between SAT-Nano, SAT-Pro and nnU-Nets on different regions and lesions. (a) Dice score comparison; (b) NSD score comparison; (c) Model size comparison. SAT-Pro is approximately 5 times smaller than the total size of all the specialist nnU-Nets, while SAT-Nano is merely one twentieth the total size of all the specialist nnU-Nets.
@@ 2.1 Main Results
Table 1: Segmentation results on different regions of human body and lesion. The best results are underlined. H&N: head and neck, LL: lower limb, UL: upper limb, WB: whole body.
@@
Metric Method Brain H&N UL Thorax Spine Abdomen LL Pelvis WB Lesion
@@
SAT-Pro 77.77 82.45 93.14 84.80 72.87 72.94 82.50 78.66 79.61 44.62
SAT-Nano 72.95 79.15 90.94 79.68 73.68 67.83 78.42 74.89 75.36 38.17
2.1.2 Class-wise Results.
For a more detailed evaluation, we show the results on all 497 classes in Fig. 4 and Fig. 5. Classes from the same region are clustered together and sorted from high to low according to the results of SAT-Pro. The scores of SAT-Pro, SAT-Nano, nnU-Nets on the same class are plotted in a vertical line. Statistics over all classes show that SAT-Pro outperforms SAT-Nano on most classes: 439/497 on DSC, and 442/497 on NSD; Meanwhile, SAT-Pro exceeds nnU-Nets on 133/497 classes on DSC, and 192/497 classes on NSD, including some important segmentation classes such as liver (+2.15 on DSC & +1.43 on NSD), pancreas (+0.83 & +2.76), lumbar vertebrae (+10.69 & +14.05). Averaging over all 497 classes, SAT-Pro achieves 78.73 on DSC, which is about 4.26% improvement over SAT-Nano (75.51); and 77.71 on NSD, about 5.31% improvement over SAT-Nano (73.76). Compared with nnU-Nets, SAT-Pro achieves 97.55% the performance of nnU-Nets on DSC and 99.32% the perfomance of nnU-Nets on NSD, demonstrating very competitive segmentation performance as an universal model. We report the detailed class-wise and dataset-wise results in Table 3, Table 4 and Table LABEL:tab:dataset_results.
Figure 4: DSC scores of SAT-Nano, SAT-Pro and nnU-Nets on each classes. Classes from the same region are clustered together and sorted from high to low according to the results of SAT-Pro.
Figure 5: NSD scores of SAT-Nano, SAT-Pro and nnU-Nets on each classes. Classes from the same region are clustered together and sorted from high to low according to the results of SAT-Pro.

### v3

@@ Abstract
Abstract. In this study, we aim to build up a model that can Segment Anything in radiology scans, driven by Text prompts, termed as SAT. Our main contributions are three folds: (i) for dataset construction, we construct the first multi-modal knowledge tree on human anatomy, including 6502 anatomical terminologies; Then we build up the largest and most comprehensive segmentation dataset for training, by collecting over 22K 3D medical image scans from 72 segmentation datasets, across 497 classes, with careful standardization on both image scans and label space; (ii) for architecture design, we propose to inject medical knowledge into a text encoder via contrastive learning, and then formulate a universal segmentation model, that can be prompted by feeding in medical terminologies in text form; (iii) As a result, we have trained SAT-Nano (110M parameters) and SAT-Pro (447M parameters), demonstrating comparable performance to 72 specialist nnU-Nets trained on each dataset/subsets. We validate SAT as a foundational segmentation model, with better generalization ability on external (unseen) datasets, and can be further improved on specific tasks after fine-tuning adaptation. Comparing with interactive segmentation model, for example, MedSAM, segmentation model prompted by text enables superior performance, scalability and robustness. As a use case, we demonstrate that SAT can act as a powerful out-of-the-box agent for large language models, enabling visual grounding in clinical procedures such as report generation. All the data, codes, and models in this work have been released.
@@ 1 Introduction
Figure 1: In contrast to conventional specialist models (a) that develop specialized solution for each task, or recently proposed interactive universal segmentation models (b) relying on real-time human interventions, Segment Anything with Text (SAT) directly takes 3D volumes as inputs, and use text as prompts to perform a wide array of medical image segmentation tasks across different modalities, anatomies, and body regions (c). It can be easily applied to clinics or seamlessly integrated with any agent-based large language model.
@@
There is a growing interest in developing universal models for medical image segmentation [54, 26], by adapting the pre-trained Segment Anything Model (SAM) [34] models from the computer vision community. However, while transferring to medical scenarios, these models trained on natural images suffer from fundamental limitations: (i) models typically perform 2D slice segmentation, that are later fused into 3D volumes through post-processing. This approach overlooks the critical contextual information inherent in 3D radiological imaging; (ii) models often require point or box inputs as prompts, thus is an interactive segmentation model, requiring considerable manual efforts for use in practice; (iii) models suffer from significant domain gaps, from image statistics to domain-specific medical knowledge.
@@
In this paper, we present the first knowledge-enhanced universal model for 3D medical volume segmentation, with language/text as prompt, termed as SAT. In practice, our model can effectively take 3D volume as visual inputs along with text prompts, to seamlessly tackle various medical image segmentation tasks, across modalities, anatomies, and body regions. As illustrated in Fig. 1, our proposed method distinguishes from previous medical segmentation paradigms, that can be seamlessly applied to clinics or integrated with any large language model. Specifically, we make the following contributions:
@@
On dataset construction, we construct a knowledge tree on anatomy concepts and definitions throughout the human body. On the visual side, we curate over 22K 3D medical image scans with 302K anatomical segmentation annotations, covering 497 categories from 72 publicly available medical segmentation datasets, termed as SAT-DS. To the best of our knowledge, SAT-DS denotes the largest and most comprehensive collection of public 3D medical segmentation datasets. To achieve such goal, we have invested significant efforts in standardizing datasets and unifying annotation labels, paving the way for universal segmentation task. For complete list of datasets and download links, we refer the reader to Table 26 and Table 4.
@@
Figure 2: Overview of our dataset for training universal segmentation model, from a wide range of segmentation tasks across various modalities and anatomical regions of the human body, including the brain, head and neck, thorax, spine, abdomen, upper limb, lower limb, and pelvis.
@@
On architecture design and training strategy, we build a universal medical segmentation model, that enables flexible segmentation across a spectrum of medical imaging modalities and anatomies, with text as prompts. Specifically, we adopt knowledge-enhanced representation learning, leveraging textual anatomical knowledge and atlas segmentation of specific anatomical structure to train the text encoder. Through this training process, the visual features of these anatomical structures are aligned with their corresponding text descriptions in the latent space, which is validated to boost the segmentation performance, especially in a long-tail distribution. Later, the text embeddings of anatomy/abnormality are treated as queries in a Transformer-based architecture, iteratively attending the visual features to extract task-specific features for precise segmentation of the queried target. To meet the requirements from different computational resources, we train two models of varying sizes, namely, SAT-Nano and SAT-Pro, and validate the effectiveness of the scaling law in universal segmentation.
@@
On experiment evaluation, we devise comprehensive metrics for universal medical segmentation across various aspects, including region-wise average, organ-wise average, and dataset-wise average. Through extensive internal and external experiments, we demonstrate that:
@@
  • Building on the unprecedented dataset collection, SAT can handle a wide range of downstream segmentation tasks with flexible text prompts, simplifying the conventional solution to train and deploy numerous specialist models. On internal evaluation, SAT-Pro, with only 447M parameters, shows comparable performance to specialist models trained individually on each dataset (72 separate nnU-Nets around 2.2B parameters); While on external evaluation, SAT-Nano, with only 110M parameters, demonstrate even better generalization ability over nnU-Net models.
  • As a foundational segmentation model pre-trained on the large-scale and comprehensive data, SAT-Pro could be further optimized and adapted to specific tasks, balancing the clinical needs between unievrsal solution and specialist models. On most of the datasets, it shows superior performance over the counterpart nnU-Nets.
  • Driven by text prompts, SAT outlines a novel universal segmentation paradigm as opposed to previous efforts in building interactive models, which relies on spatial prompts. This could save tremendous manual efforts from prompting in clinical applications. On performance, we demonstrate it can outperform the state-of-the-art interactive universal segmentation model, namely, MedSAM, on internal and external evaluations, while being robust to targets with ambiguous spatial relationships.
  • The text-prompted feature and universality of makes SAT a powerful out-of-box agents for language model. We show SAT can be seamlessly integrated with any large language models such as GPT-4 [2], automatically providing explainable grounded report in clinical scenarios, or providing grounding annotation for medical data in a scalable way.
@@ 2 Results
The objective of our proposed model, termed Segment Anything with Text (SAT), is to develop a universal segmentation framework for 3D medical images, guided by text prompts, to handle various segmentation tasks using a single model. The universality of SAT not only allows easy adaptation to clinical procedures, meeting a wide range of clinical demands, but also makes SAT a versatile tool that bridges linguistic inputs and segmentation outputs. To this end, we have developed a large-scale and comprehensive data collection, termed SAT-DS, covering a broad spectrum of medical image segmentation tasks. It includes 497 anatomical targets across 8 regions and various lesions of the human body, assembled from 72 distinct datasets.
In this section, we will detail the experiment results where SAT is comprehensively evaluated against two representative baselines: nnU-Nets [27], an ensemble of specialist models optimizing configuration and training individually on each dataset; and MedSAM [54], state-of-the-art interactive universal segmentation model driven by box prompts. The evaluation results are present from multiple perspectives, including region-wise, class-wise, and dataset-wise, to gain a deep understanding of our models’ capabilities. Note that class-wise and region-wise evaluations involved averaging the results from different nnU-Nets. For instance, the performance metrics for the ‘brainstem’ in CT images represent the macro average from models trained on datasets like ‘HAN Seg’, ‘PDDCA’, and ‘SegRap2023 Task 1’, all of which include annotations for this anatomical class. Detailed experiment settings can be found in Sec. 4.4.
2.1 Comparison with Specialist Models on Automatic Segmentation
Evaluation on Internal Datasets. In this experiment setting, we aim to evaluate SAT on in-domain samples, i.e.,3D images that follow the same distribution as train set. Specifically, we split each of the 72 datasets in SAT-DS into train and test splits in 8:2 ratio. We train and evaluate two variants SAT-Pro and SAT-Nano in this setting; While for nnU-Nets, we train 72 nnU-Nets specialized on each dataset and report their results as a whole. Table 1 shows the region-wise results on 8 regions of human body, including Brain, Head and Neck, Thorax, Abdomen, Pelvis, Spine, Upper Limb, and Lower Limb, as well as Lesion, in terms of DSC and NSD, respectively. Classes existing in multiple regions are specifically grouped as Whole Body. Notably, SAT-Pro consistently outperforms nnU-Nets in four regions: on Head and Neck, it surpasses nnU-Nets by 5.21 on DSC and 8.27 on NSD; On Thorax, it surpass nnU-Nets by 0.78 on DSC and 2.49 on NSD; On Upper Limb, it surpasses nnU-Nets by 2.8 on DSC and 9.28 on NSD; On Lower Limb, it surpasses nnU-Nets by 1.53 on DSC and 2.48 on NSD. Fig. 3 provides another view on the above results, where it can be seen that SAT-Pro shows slightly lower or comparable segmentation performance to the 72 nnU-Nets, while being of significantly smaller size, approximately 1/5 of the assemble of nnU-Nets. We present more detailed region-wise results in supplementary material Fig. 13 and Fig. 14. We further present class-wise results in supplementary material Sec. 8.3, and dataset-wise results in supplementary material Sec. 8.2.
@@ 2.1 Comparison with Specialist Models on Automatic Segmentation
Table 1: Internal evaluation results of nnU-Nets, SAT-Nano, SAT-Pro and SAT-Pro-FT on different regions of human body and lesions. H&N: head and neck, LL: lower limb, UL: upper limb, WB: whole body, All: average over all the 497 classes. The best results are bolded.
@@
Metric Method Brain H&N UL Thorax Spine Abdomen LL Pelvis WB Lesion All
@@
SAT-Nano 72.95 79.15 90.94 79.68 73.68 67.83 78.42 74.89 75.36 38.17 73.76
SAT-Pro 77.77 82.45 93.56 85.15 72.87 72.94 82.50 78.66 79.61 44.87 77.82
SAT-Pro-Ft 80.43 84.01 95.64 86.38 82.37 78.42 87.25 82.92 82.48 47.74 81.60
Figure 3: Comparison between SAT-Nano, SAT-Pro and nnU-Nets on different regions and lesions. (a) DSC score comparison; (b) NSD score comparison; (c) Model size comparison. SAT-Pro is approximately 5 times smaller than the total size of all the specialist nnU-Nets, while SAT-Nano is merely one twentieth the total size of all the specialist nnU-Nets.
Dataset-specific Adaptation. Built on the unprecedented data collection, SAT can be used as a strong pre-trained model, and adapted towards specific downstream dataset, by further finetuning on the training split of certain dataset. As shown in Table 1, we also report the region-wise results of SAT-Pro-FT. When comparing to the original SAT-Pro, it shows significant improvement on all the regions and lesion classes, especially on Spine (+7.75 on DSC & +9.5 on NSD) and Pelvis (+3.83 on DSC & +4.26 on NSD). Overall, after finetuning on each of the 72 datasets, and compute the average, SAT-Pro-Ft improves 2.6 on DSC and 3.78 on NSD, ultimately outperforms nnU-Nets by 0.87 on DSC, and 3.45 on NSD. The class-wise results are presented in supplementary material Sec. 8.3, and dataset-wise results in supplementary material Sec. 8.2.
Evaluation on External Datasets. In this experiment setting, we aim to evaluate the generalization performance segmentation models on images outside the train set distribution, e.g.,to data from different centers. Note that, this do not involve new classes, as the segmentation targets in human body are relatively limited and fixed. To this end, we use a subset collection of SAT-DS as training data, termed as SAT-DS-Nano, covering 49 datasets, and 429 classes; and take 10 held-out datasets for evaluation. The details of SAT-DS-Nano can be found in Sec. 4.4, and the composition can be found in Table 26. Though we found Although SAT-Pro is generally superior over SAT-Nano in internal evaluation, to save computational cost, we only train and evaluate SAT-Nano in this setting; While to systematically evaluate nnU-Nets, for each held-out dataset, we traverse all the 49 datasets and corresponding models, and evaluate them on all the classes shared between the two datasets. For example, to segment ‘lung’ in COVID-19 CT Seg, we utilize nnU-Nets trained on CT-ORG and LUNA16, as they all involve this class, and then average the results. Overall, 23 out of 49 nnU-Nets are transferred for 80 times and are evaluated as a whole. The transfer details are in supplementary material Fig. 19. To maintain nnU-Nets’ performance, the target dataset are processed consistently with the source dataset in transfer.
Table 2: External evaluation results on each held-out dataset. AA stands for AbdomenAtlasMini1.0 [71] (a very recently released segmentation dataset not included in SAT-DS); CHAOS stands for CHAOS CT; DA stands for DAP Atlas; Liver stands for MSD Liver; Pancreas stands for MSD Pancreas; Prostate stands for MSD Prostate; Spleen stands for MSD Spleen; TS V2 stands for TotalSegmentator V2. Here, SAT-Nano denotes the variant trained with held-out datasets. The best results are bolded. Note that, nnU-Nets and SAT-Nano are fully automatic methods, while MedSAM is interactive. Datasets involved in training MedSAM are excluded and marked /.
Metric Method AA BTCV CHAOS COVID19 DA Liver Pancreas Prostate Spleen TS V2
[…]
SAT-Nano 83.56 76.01 96.69 94.6 61.53 96.58 87.21 72.76 93.68 46.8
Oracle Box 75.26 68.82 72.93 / 54.52 73.13 / / / /
MedSAM (Tight) 88.41 77.94 96.48 / 64.03 95.87 / / / /
MedSAM (Loose) 75.39 56.01 94.15 / 41.07 48.66 / / / /
[…]
SAT-Nano 75.59 71.94 82.35 86.83 57.65 79.46 75.62 44.28 81.71 37.76
Oracle Box 45.92 54.36 28.28 / 44.64 26.31 / / / /
MedSAM (Tight) 78.43 74.4 81.0 / 55.03 77.26 / / / /
MedSAM (Loose) 45.7 36.04 66.87 / 21.47 27.89 / / / /
We report dataset-wise results in Table 2. In general, SAT-Nano outperforms nnU-Nets on 9 out of 10 datasets on DSC score (+2.21), and excels consistently on all datasets on NSD score (+6.73). Notably, SAT-Nano exceeds nnU-Nets significantly on datasets like MSD Prancreas (+13.94 on DSC & +22.99 on NSD), MSD Liver (+1.17 on DSC & +10.82 on NSD) and MSD Prostate (+8.91 on DSC & +1.72 on NSD). We conjecture this is primarily because the dataset-wise customization hinders the generalization of nnU-Nets to external datasets. Note that, based on the universal segmentation capability, SAT-Nano only needs the user to specify the name of the interested region at inference time, while nnU-Nets require cumbersome transfers from models specialized on different datasets and class sets (label space).
2.2 Comparison with Interactive Segmentation Model
In this section, we compare our proposed SAT with MedSAM, As MedSAM is an out-of-the-box universal segmentation methods trained on large-scale data, we directly use it on our data without finetuning. Similar to Sec. 2.1, the comparison with MedSAM is conducted in both internal evaluation and external evalution. However, due to the inconsistent training data, we carefully choose 32 out of 72 datasets for internal evaluation, which are also involved in training MedSAM. Note that, even though these datasets are included in training MedSAM, the train-test splits are not aligned, which means the test set we used might be leaked in training MedSAM. The detailed composition can be found in supplementary material Table 10. And 5 out of 10 held-out datasets are selected for external evaluation, which are excluded in training MedSAM. In both internal and external evaluation, SAT-Pro and SAT-Nano refer to the same models in Sec. 2.1, only the number of evaluation datasets are reduced. We report three results for MedSAM: simulate box prompts based on the ground truth segmentation, with minimum rectangle covering the ground truth (denoted as Tight), i.e., the most accurate prompts; randomly shift each box corner up to 8% of the image resolution (denoted as Loose), i.e., allow errors to some extent; directly use the tight box prompts as prediction (denoted as Oracle Box), i.e., the input baseline of MedSAM.
Evaluation on Internal Datasets. Table 3 shows the region-wise results of all methods. Notably, SAT-Pro consistently outperforms MedSAM on all regions in terms of both DSC ($+7.08$) and NSD ($+10.8$), even if MedSAM is prompted with the most accurate box (Tight). On lesion segmentation, SAT-Pro underperforms MedSAM (Tight), due to the small size of lesion, and the box prompts to MedSAM have already provided very strong prior, as indicated by the results of inputting oracle box, which even shows better performance than MedSAM output on DSC score. We present more detailed region-wise results in supplementary material Fig. 17 and Fig. 18, class-wise results in supplementary material Sec. 8.6 and dataset-wise results in supplementary material Sec. 8.5.
Table 3: Internal evaluation results of SAT-Pro, SAT-Nano and MedSAMs on different regions of human body and lesion. H&N: head and neck, All: average over all the classes. The best results are bolded. Note that SAT-Pro and SAT-Nano are fully automatic methods, while MedSAM is interactive.
Metric Method Brain H&N Thorax Spine Abdomen Limb Pelvis Lesion All
[…]
SAT-Nano 69.05 77.93 79.06 81.88 78.43 77.6 78.31 47.7 76.66
Oracle Box 55.84 85.32 72.85 78.11 71.92 78.96 81.3 67.94 63.57
MedSAM (Tight) 54.35 78.48 73.01 79.09 77.35 80.53 84.42 65.85 75.39
MedSAM (Loose) 14.7 44.68 36.49 45.72 19.79 42.97 51.42 15.71 35.18
[…]
SAT-Nano 75.62 79.44 77.74 82.7 74.96 77.58 72.98 40.38 74.82
Oracle Box 60.55 86.83 66.16 69.86 59.51 70.67 71.59 61.59 48.93
MedSAM (Tight) 68.62 82.37 68.26 73.29 73.39 73.84 74.9 62.28 70.99
MedSAM (Loose) 10.29 43.57 31.61 43.84 16.81 41.54 46.74 11.46 31.50
We further present the qualitative results from two representative examples to illustrate the performance gap between SAT-Pro with MedSAM. In Fig. 4, the upper row shows myocardium segmentations on a CT image. As the myocardium is hard to be referred by the rectangular box prompt, MedSAM generate unsatisfactory prediction which wrongly includes the left heart ventricle surrounded by the myocardium. By comparison, SAT-Pro generate the desirable prediction via the word ‘myocardium‘. It reveals the fundamental dilemma underlying MedSAM: spatial prompts can be ambiguous in complicated situations. The lower row demonstrates colon cancer segmentations on a CT image. The tight box prompt MedSAM receives forms an acceptable segmentation, despite its limitation as a rectangle. While by contrast, the segmentation output from MedSAM is even worse. This confirms our conjecture that MedSAM performs well in lesion targets primarily due to the strong prior provided by the tight box prompt.
Figure 4: Visualization of MedSAM (Tight) and SAT-Pro results on liver (upper row) and colon cancer (lower row) segmentation examples. The ground truth and segmentation masks are painted in red, while box prompts of MedSAM are plotted in blue. The DSC score is calculated in slice-wise manner.
Evaluation on External Datasets. The dataset-wise results are shown in Table 2. Compared with MedSAM (Tight), SAT-Nano achieves higher DSC score on CHAOS CT (+0.21) and MSD Liver (+0.71); and higher NSD score on CHAOS CT (+1.35), DAP Atlas (+2.62) and MSD Liver (+1.8). While for MedSAM (Loose), the performance drops significantly on all datasets, indicating that the model is sensitive to user inputs.